OpenStudy (anonymous):

In the square JKLM below, the length of slide KL is 66 ft What is the perimeter?

OpenStudy (anonymous):

given the length of the square be a units then perimeter =4*a here a=66

OpenStudy (anonymous):

so just multiply 4 and 66

OpenStudy (anonymous):

I don't he would be asking if the question was that easy. The question is really asking that if the length of KL is 66 feet, and KL is the diagonal of the square, then what is the perimeter. Let's from a diagram. Since the diagonal creates a right triangle with 2 equal sides, we can apply the Pythagorean Theorem here: a^2 + b^2 = c^2 (x)^2 + (x)^2 = 64^2 2x^2 = 4096 x^2 = 2048 x = 45.253 Therefore, x, which is the side length of the square, is approximately 45.253. Since the Perimeter = 4x, then P = 4(45.253) = 181.012 feet

OpenStudy (anonymous):

@MandyMicheala